OpenStudy (anonymous):

(a)find y as a function of x, if y'=e^(x-1) and y =1 when x=1. What is the y intercept of this curve? (b)The gradient of a curve is given by y'=e^(2-x) and the curve passes through the point (0,1). What is the equation of this curve? What is its horizontal axis? (c)f=Find y, given that y'=2^(-x), and y=1/2log2 when x =0. (d) It is known that f'(x) = e^x + 1/e and that f(-1)=-1. Find f(0)

OpenStudy (anonymous):

\[\int\limits e ^{x-1}\]

OpenStudy (anonymous):

\[1/e \int\limits e^x d x\]

OpenStudy (dumbcow):

a) \[y = \int\limits_{}^{} e ^{x-1}dx\] u = x-1 du = dx \[y=\int\limits_{}^{}e ^{u}du = e ^{u}+C=e ^{x-1}+C\] 1=e^0 +C = 1+C C=0 y = e^x-1

OpenStudy (anonymous):

\[1/e[e^x +c]\]

OpenStudy (anonymous):

i actually did (a) i need help with the other ones more lol

OpenStudy (dumbcow):

b) y = -e^2-x + C 1=-e^2 +C C = 1+e^2 \[y = -e ^{2-x} +e ^{2}+1\] c) y=-2^(-x)/ln2 +C 1/2ln2 = -1/ln2 +C C = 3/2ln2 \[y = -\frac{1}{\ln 2}(2^{-x} -\frac{3}{2})\] d) f(x) = e^x + x/e +C -1 = e^-1 -1/e +C C = -1 f(x) =e^x +x/e -1 f(0) = 1-1 = 0

OpenStudy (anonymous):

for c is it the same as writing y=1-2^(-x)/log2

OpenStudy (dumbcow):

yeah well i get 3/2 or 1.5 instead of 1 but yes you could write it either way

OpenStudy (anonymous):

thats the answer at the back of the book lol

OpenStudy (dumbcow):

ahh well book must be wrong..:) is it 1/2*ln2 or 1/(2ln2) for initial conditions

OpenStudy (anonymous):

isnt it the same thing

OpenStudy (dumbcow):

no one the log is on top the other its on the bottom

OpenStudy (anonymous):

oh its 1/(2log2)

OpenStudy (dumbcow):

thats what i thought i dont know what i did wrong on that one...

OpenStudy (anonymous):

when i did it i got -c=-1/2log2 -1/2log2 c=1/2log2+1/2log2

OpenStudy (anonymous):

theres should be no negative infront of y=-2^(-x)/ln2 +C i think lol

OpenStudy (dumbcow):

wait is x=0 or x=1 ? no there needs to be a neg because derivative of -x = -1

OpenStudy (anonymous):

x = 1

OpenStudy (anonymous):

oh sorry i wrote it wrong

OpenStudy (dumbcow):

thats it then lol it says 0 above

OpenStudy (anonymous):

yeah im sorry lol i must of did it by accident lol

OpenStudy (dumbcow):

it happens, keeps me on my toes :)

OpenStudy (anonymous):

lol thanks heaps you made me learn something new today:)

OpenStudy (dumbcow):

your welcome good luck on the rest

OpenStudy (anonymous):

for question (b) when you sub in x = -1 dont you just get -e

OpenStudy (dumbcow):

f(x) = e^x + x/e +C when x=-1, = e^-1 + -1/e +C =1/e - 1/e +C =0 + C its not -e because e is in denominator

OpenStudy (dumbcow):

oh i was looking at d)...is that right?

OpenStudy (anonymous):

no (b) lol

OpenStudy (anonymous):

ur answer is right but i tried doing it on my own and got -e :(

OpenStudy (dumbcow):

ok y=−e^(2−x)+e^2+1 x=-1 =-e^(2-(-1)) +e^2 +1 =-e^3 +e^2 +1 ahh no you cant combine the e's because of the addition

OpenStudy (dumbcow):

oh why are you using x=-1

OpenStudy (anonymous):

omg whats wrong with me lol im not reading any of the questions right im really sorry for troubling you for no reason

OpenStudy (dumbcow):

haha no prob